# EchoONE: Segmenting Multiple echocardiography Planes in One Model

### v2

@@ 4.3.1 Internal evaluation
Robustness to multiple planes: We first shown the results for internal evaluation in Table 2 in an plane-wise manner to better evaluate the robustness of these methods for MPS problem. For each plane, the results from the five internal datasets are averaged. Among the four traditional models, U-Net, SwinUNet, and H2Former generally perform better on our dataset compared to most of the SAM-based models, with the exception of SAMUS, which is specifically designed for ultrasound image segmentation. The rest SAM-based methods, although retrained using the same data, only achieve reasonable good performance, and show clear lower performance than EchoONE for all the planes. It’s worth to note that EchoONE performs significantly better than all competitors for the PSAX plane, which may be attributed to the uniform design of the semantice-aware PC-Mask for dense prompt.
Robustness across different structures: Figure 6 presents a more detailed Dice (%) score for each cardiac structure (LV, LA, MYO) of different planes. It can been drawn that our method consistently achieves higher Dice score than all the competitors for all these cardiac structures.
Robustness across datasets: We also display the dataset-wise results in Table 3 for A, B, C, D, and Figure 7 for CAMUS. Average is taken across planes of each dataset. The proposed EchoONE outperforms all the competitors in all five internal datasets.
@@ 4.3.2 External evaluation
Results for external evaluation is shown in Table 4. Without training with the two datasets, our method performances the best for both of them, revealing its great generalization. Even for HMC_QU, where low quality images exists and annotations are noisy, our method still deliver a Dice score of 73.94%. These observations demonstrate the great potential of EchoONE for images of real clinical practice.
@@
Table 3: Dataset-wise results of EhoONE and other models on the evaluation set of the four private internal datasets.
@@ 4.4 Ablation studies
To further investigate the effectiveness of our dense prompt generation module PC-Mask and the fusion and adaption module (LFFA), ablation studies are conducted on the test data of our five internal datasets. Results are shown in Table 5.
Effectiveness of PC-Mask: When only PC-Mask is introduced, our method can improve the segmentation performance for all planes, with gain of 0.06%, 0.54%, 1.01% and 1.48%, respectively. Especially for the PSAX, which differs most from the rest three planes, PC-Mask can effectively embed the prior structure through the dense prompt, thereby leading to obvious improvement from 74.86% to 76.34% in term of mIoU. When LFFA is included, PC-Mask can still bring further improvements over all planes.
@@
From these observations, it can be drawn that By leveraging both the prior information in a semantic-aware way and the powerful SAM model with lcoal features fusion and adaption, EchoONE can achieve accurate and robust performance for the MPS problem of echocardiographic images.
@@ 5 Conclusion
In this paper, we propose EchoONE, a model designed to address the challenging multi-plane segmentation problem in echocardiography images. We propose a novel dense prompt learning module PC-Mask by leveraging the prior structure knowledge in a composable way, therefore provides effective plane-specific semantic guidance during segmentation. We also introduce a learnable CNN local feature branch to tune the image encoder and adapt the mask decoder. A newly-designed LLFA module is introduced for the fusion, which not only enhances the final performance, but also accelerates the convergence. This is the fist effective solution aiming to segment all ehocardiogrpahic planes with one uniform model. This will definitely simplify the deployment of AI techniques in clinical practice. Although only validated with ultrasound images, our method is also generalizable to MPS problmes in other medical image modalities.

### v3

@@ 4.3.1 Internal evaluation
Robustness to multiple planes: We first show the results for internal evaluation in Table 2 in an plane-wise manner to better evaluate the robustness of these methods for MPS problem. For each plane, the results from the five internal datasets are averaged. Among the four traditional models, U-Net, SwinUNet, and H2Former generally perform better on our dataset compared to most of the SAM-based models, with the exception of SAMUS, which is specifically designed for ultrasound image segmentation. The rest SAM-based methods, although retrained using the same data, only achieve reasonable good performance, and show clear lower performance than EchoONE for all the planes. It’s worth to note that EchoONE performs significantly better than all competitors for the PSAX plane, which may be attributed to the uniform design of the semantic-aware PC-Mask for dense prompt.
Robustness across different structures: Figure 6 presents a more detailed Dice (%) score for each cardiac structure (LV, LA, MYO) of different planes. It can be drawn that our method consistently achieves a higher Dice score than all competitors for all these cardiac structures.
Robustness across datasets: We also display the dataset-wise results in Table 3 for A, B, C, D, and Figure 7 for CAMUS. Average is taken across planes of each dataset. The proposed EchoONE outperforms all competitors in all five internal datasets.
@@ 4.3.2 External evaluation
Results for external evaluation are shown in Table 4. Without training with the two datasets, our method performs the best for both of them, revealing its great generalization. Even for HMC_QU, where low quality images exist and annotations are noisy, our method still delivers a Dice score of 73.94%. These observations demonstrate the great potential of EchoONE for images of real clinical practice.
@@
Table 3: Dataset-wise results of EchoONE and other models on the evaluation set of the four private internal datasets.
@@ 4.4 Ablation studies
To further investigate the effectiveness of our dense prompt generation module PC-Mask and the fusion and adaption module (LFFA), ablation studies are conducted on the test data of our five internal datasets. The results are shown in Table 5.
Effectiveness of PC-Mask: When only PC-Mask is introduced, our method can improve the segmentation performance for all planes, with mIoU gains of 0.06%, 0.54%, 1.01% and 1.48%, respectively. Especially for PSAX, which differs most from the other three planes, PC-Mask can effectively embed the prior structure through the dense prompt, thereby leading to obvious improvement from 74.86% to 76.34% in terms of mIoU. When LFFA is included, PC-Mask can still bring further improvements on all planes.
@@
From these observations, it can be drawn that By leveraging both the prior information in a semantic-aware way and the powerful SAM model with local features fusion and adaption, EchoONE can achieve accurate and robust performance for the MPS problem of echocardiographic images.
@@ 5 Conclusion
In this paper, we propose EchoONE, a model designed to address the challenging multi-plane segmentation problem in echocardiography images. We propose a novel dense prompt learning module PC-Mask by leveraging the prior structure knowledge in a composable way, therefore provides effective plane-specific semantic guidance during segmentation. We also introduce a learnable CNN local feature branch to tune the image encoder and adapt the mask decoder. A newly-designed LFFA module is introduced for the fusion, which not only enhances the final performance, but also accelerates the convergence. This is the first effective solution that aims to segment all echocardiographic planes with one uniform model. This will definitely simplify the deployment of AI techniques in clinical practice. Although validated only with ultrasound images, our method is also generalizable to MPS problems in other medical image modalities.
6 Acknowledgment
The work is partially supported by NSFC (62471313, 12326619, 62306183, 62171290), NSF of Guangdong (2024A1515030143, 2024A1515010194, 2023A0505020002), Shenzhen Science and Technology Program (JCYJ20240813141807010), Frontier Technology Development Program of Jiangsu (No.BF2024078), CAMS Innovation Fund for Medical Sciences (CIFMS) (2023-I2M-C&T-B-117, 2023-I2M-C&T-B-056), and internal project of National Engineering Laboratory for BDSC (SZU-BDSC-IF2024-02).